OpenStudy (anonymous):

Find the average rate of change of f(x) = x^3 - 3x +3 from -9 to -5

OpenStudy (mertsj):

Find the f(-5) Find the f(-9) Then do this calculation: \[\frac{f(-5)-f(-9)}{-5-(-9)}\]

OpenStudy (anonymous):

so just -5 - -9/-5 -9?

OpenStudy (mertsj):

What is the function f?

OpenStudy (anonymous):

-14?

OpenStudy (mertsj):

f(x) = x^3 - 3x +3 That is the function f.

OpenStudy (mertsj):

Replace x with -5

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

63

OpenStudy (anonymous):

Which is prob wrong

OpenStudy (mertsj):

\[(-5)^3-3(-5)+3=-125+15+3=-110+3=-107\]

OpenStudy (mertsj):

Now you find the f(-9)

OpenStudy (anonymous):

Yea you just did a lot that I dont understand ;(

OpenStudy (anonymous):

Sorry math is my weak sub

OpenStudy (mertsj):

Well, I would like to help you but I don't know where to start. do you know how to raise a number to a power?